OpenStudy (anonymous):

suppose z=f(x,y) is express in polar form z=(r, @) by making x=rcos@ and y=rsin@

OpenStudy (anonymous):

couldnt open question at first

OpenStudy (anonymous):

so dx/dr = cos @ what is next

hartnn (hartnn):

u need dr/dx, right ? dr/dx = 1/ (dx/dr)

OpenStudy (anonymous):

you mean 1=cos @ dr/dx?

hartnn (hartnn):

yeah, so dr/dx = 1/cos@

OpenStudy (anonymous):

but the answer said it is cos @ and not 1/cos@

hartnn (hartnn):

ok, let me try

hartnn (hartnn):

r^2 = x^2+y^2

hartnn (hartnn):

2r dr/dx = 2x dr/dx = x/r

hartnn (hartnn):

dr/dx = rcos @/r

hartnn (hartnn):

hmm....i am getting cos @ now...

hartnn (hartnn):

did u get what i did ?

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

the wording of the question is just confusing

OpenStudy (anonymous):

thanks again

hartnn (hartnn):

welcome ^_^